A SKYDIVING CLUB HAS 30 MEMBERS.THE CLUB MEMBERS JUMP IN PAIRS.HOW MANY DIFFERENT PAIRS OF SKYDIVERS ARE POSSIBLE.
Answer by stanbon(75887)

A SKYDIVING CLUB HAS 30 MEMBERS.THE CLUB MEMBERS JUMP IN PAIRS.HOW MANY DIFFERENT PAIRS OF SKYDIVERS ARE POSSIBLE.
----
Ans: 30C2 = (30*29)/(1*2) = 15*29 = 425 pairs
Cheers,
Stan H.